OpenStudy (anonymous):

find all solutions to the equation 2 cos0 +1=0

OpenStudy (across):

\[2\cos\theta+1=0.\]Is this what you mean?

OpenStudy (anonymous):

yeass, i couldnt do the zero thingy

OpenStudy (across):

Well, let's do this by steps: 1. You're given: \(2\cos\theta+1=0\). 2. Our goal is to solve (1) for \(x\): 3. \(2\cos\theta=-1\), 4. \(\cos\theta=-1/2\), 5. \(\theta=\cos^{-1}(-1/2)=2\pi/3\)\[\]

OpenStudy (across):

Since \(\cos\theta\) is a periodic function with period \(2\pi\), our solutions will take the form \(\theta=2\pi n-2/3\pi\) and \(\theta=2\pi n+2/3\pi\) where \(n\in\mathbb{Z}\).\[\]